OpenStudy (anonymous):

plz solve this question and answer it soon i need the answer quickly plz : the question is : if tan35=2a-1 then solve (sin145-sin235)/cos325

myininaya (myininaya):

2a=tan35+1 a=1/2*(tan35+1)

OpenStudy (anonymous):

0.85010376910485488972926135972242

OpenStudy (anonymous):

dear duke we dont neet to find a we need to solve (sin145-sin235)/cos325

myininaya (myininaya):

there is nothing to solve there we solve equations that is an expression do you mean approximate or simplify?

OpenStudy (anonymous):

translation maybe but one statement stand on it's on as myininaya states

OpenStudy (anonymous):

approximate

myininaya (myininaya):

1.700207538

OpenStudy (anonymous):

1.7002075382097097794585227194448 same as myininaya

OpenStudy (anonymous):

myninaya can u write the way u solved plz?

myininaya (myininaya):

but with a closer approximation lol

myininaya (myininaya):

i mean u do duke have the closer one

myininaya (myininaya):

ok roh do you have a calculator

OpenStudy (anonymous):

oh what type of calculator? hp or ti order of operations is the problem most likely

myininaya (myininaya):

go to mode put in degrees

OpenStudy (anonymous):

lolz any way without calculator through any trignmoetry formula

OpenStudy (anonymous):

((sin(145))-(sin(235)))/(cos(325)) in ti32

OpenStudy (anonymous):

i did like this : (sin(90+55)-sin(90+55))/cos(270+55)

myininaya (myininaya):

sin(145)-sin(235) =sin(145)-sin(145+90) =sin(145)-[sin(145)cos(90)+sin(90)cos(145)] =sin(145)-[sin(145)*0+1*cos(145)] =sin(145)-cos(145) --------------------- cos(325)=cos(180+145)=cos(145)cos(180)-sin(145)sin(180) =cos(145)*(-1)-sin(145)*0 =-cos(145) so we have \[\frac{\sin(145)-\cos(145)}{-\cos(145)}=\frac{\sin(145)}{-\cos(145)}+1=-\tan(145)+1\] hmm... thinking....

myininaya (myininaya):

i dont see how to not do it with a calculator anyone else?

myininaya (myininaya):

without*

OpenStudy (anonymous):

sin (145) = sin (180- 35) = sin 35 cos(325) = cos(360 - 35) = cos35 if only there was some way to convert the sin(235) to either cos or sin of 35

OpenStudy (anonymous):

sin (270-35) = sin 235 thanks joe

myininaya (myininaya):

sin(235)=sin(270-35) lol

myininaya (myininaya):

i was gonna say that duke

OpenStudy (anonymous):

LOL im tired as hell, thats it.

myininaya (myininaya):

lol duke is awesome :)

OpenStudy (anonymous):

i agree the answer is simple now lol

OpenStudy (anonymous):

so the answer is: \[\frac{\sin(35)+\cos(35)}{\cos(35)} = \tan(35)+1 = 2a-1+1 = 2a\] ? im walking into the middle of this, trying to catch up.

OpenStudy (anonymous):

yeah i got the answer thanks all joe is right so we will have the answer like this : sin 35+sin35/ cos 35 = tan(35)+1 so we have from the question that tan 35=2a-1 so we have: 2a-1+1= 2a

OpenStudy (anonymous):

thank you joemath and myininaya

OpenStudy (anonymous):

Dont forget duke!